OpenStudy (anonymous):

Find four solutions of the equation. Write the solutions as ordered pairs. y= -3x-6. i need help fast!! im timed on this question

OpenStudy (radar):

They want four, I show you two. y=-3x-6 Let x = 0, then y= 0 - 6 = -6 ordered pair (0,-6)

OpenStudy (radar):

another let x = -1 y = (-3)(-1) - 6 = 3 -6 = -3 (-1,-3) Get the drift?

OpenStudy (radar):

Try x = 1

OpenStudy (radar):

Or x=2 would be a good one.

OpenStudy (radar):

Just plug in a chosen x value and solve for y.

OpenStudy (radar):

I know it is hard to choose, when you have a infinite number of choices.

OpenStudy (anonymous):

yea, i really dont get this, im stuck...:(

OpenStudy (radar):

O.K. choose x=2 and substitute 2 for x in your equation y = -3x -6. show your work so I can help.

OpenStudy (anonymous):

y=-3(2)-6 3*2=6 6-6=0 y= (0,6)?

OpenStudy (radar):

Your 2nd line you omitted the - sign -3*2 = -6 your 3rd line will now look like this -6-6= -12 the ordered pair would be (2,-12)

OpenStudy (radar):

Do the last one using your choice for x, suggest x=-2 Note a ordered pair, the first one is x and the 2nd one is y (x,y)

OpenStudy (anonymous):

i still dont get it cause my answer choices are a- (-4,3), (-7,15), (5,-24), (-6,12) b- (-4,6), (-7,15). (5,-21), (-6,12) c- (-4,6), (-6,15), (5,-21), (-7,12) d- (4,6), (-7,15), (5,3), (-4,12)

OpenStudy (radar):

Well well! They aren't going to let you choose, they have made the choice for you, you only have to pick out the correct one.

OpenStudy (radar):

Looking at their choices, I would test that ordered pair (-7,15) as it appears in three of the four choices. 15=-3(-7) - 6 15= 21-6 = 15 That works we just eliminated choice c. So now we are only concerned with a, b, and d. Look at the first ordered pair in choice a, the (-4,3) Is it correct?

OpenStudy (anonymous):

no, it says a is wrong

OpenStudy (radar):

3 = -3(-4) - 6 3 = 12 -6 3=6 Uh-Oh Thats wrong so a. aint it. You now have only b and d to check. check that ordered pair (5,3) in d I just bet it might be incorrec, if so then

OpenStudy (radar):

I didn't say to check number a if it is wrong, I said to check out the ordered pair (-4,3) to see if it was correct, as I show above it was incorrec so that makes choice a incorrect!.

OpenStudy (radar):

Now focus, check out the order pair (5,3) in d. to see if it is correct. The ordered pair is saying y=3 and x=5. Is that correct?

OpenStudy (anonymous):

yes

OpenStudy (radar):

Quit guessing.......work it out y=-3x-6 checking the 5,3 3= -3(5) -6 3 = -15 - 6 3 = -21 that is not true in Missouri, choice d stinks. Sorry but I gotta go.